OpenStudy (anonymous):

if h,s, V be the height, Curved surface area and volume of a cone respectively then 3piVh^3 - s^2h^2 + 9V^2 is: NOTE PI = 22/7 or 3.14

OpenStudy (anonymous):

sstarica do u want options

OpenStudy (anonymous):

you mean \[(3\pi Vh)^3 - s^2h^2 + 9V^2\] so what's the problem?

OpenStudy (anonymous):

no sstarica

OpenStudy (anonymous):

it is only h^3

OpenStudy (anonymous):

so you have to solve it

OpenStudy (anonymous):

then why do you have 3 piV multiplied by h?

OpenStudy (anonymous):

i dodn't undeand rst

OpenStudy (anonymous):

you have 3piVh ^3 right?

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

you want to solve the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but your question is unfortunately unclear, what is excatly that you want?

OpenStudy (anonymous):

SEE i am giving you V= 1/3 * pi * r^2 * h

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

AND S = pi^2r^2(root of (r^2 + h^2))

OpenStudy (anonymous):

so you don't know the answer of only 8th standard

OpenStudy (anonymous):

no now, I got you. :)

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

ANSWER IT

OpenStudy (anonymous):

excuse me, you're the one who asked for help, why are you being so boiled up? It's rude, please be polite.

OpenStudy (anonymous):

okey

OpenStudy (anonymous):

all you have to do is substitute V and S in the equation and simplify.

OpenStudy (anonymous):

THIS IS THE PROBLEM THAT I M FACING THE ANSWER IS NOT IN OPTIONS

OpenStudy (anonymous):

and when you do that you get in the end.\[\pi^2r^2h^4 - 4\pi r^4h^2(r^2 + h^2) + 1/3 \pi r^2h\]

OpenStudy (anonymous):

yes i did that

OpenStudy (anonymous):

ofcourse you won't get it corret in the first place, try again. What are the options, tell me?

OpenStudy (anonymous):

okey i m giving the options

OpenStudy (anonymous):

1) option - 2

OpenStudy (anonymous):

2) option - 0

OpenStudy (anonymous):

3) option - 2pi

OpenStudy (anonymous):

4) option - 3(pi)^2

OpenStudy (anonymous):

I've made a mistake, in the end of the answer I gave you above it's supposed to be \[3\pi r^2h\] and note 1/3 pir^2h

OpenStudy (anonymous):

not*

OpenStudy (anonymous):

the answer is 0 !

OpenStudy (anonymous):

here the solution is : first you'll substitute the given in the formula and you will end up with this: \[\pi^2 r^2h^4 - 4 \pi^2r^4h^2(r^2 + h^2) + \pi^2r^4h^2\] then simplify, it's a long step but in the end you will end up with : \[-3\pi^2r^2h^4 - 3\pi^2r^2h^2\] = \[-3\pi^2h^2(r^2h^2 - r^2h^2)\] = \[-3\pi^2h^2\] x 0 = 0 That's it :)

OpenStudy (anonymous):

wait that's wrong >_<

OpenStudy (anonymous):

this has caused me a headache._.